OpenStudy (anonymous):

Solve the DE: If G = D(xD-1) and H = (xD-1)(xD+2) Determine GH and HG.

OpenStudy (anonymous):

@hartnn

OpenStudy (perl):

what does D stand for in this context. the derivative ?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

D stands for Differential Operator

OpenStudy (perl):

i might need some theory, maybe you could post me the name of your book or something

OpenStudy (anonymous):

we don't have a book.. just handouts... the topic is about Differential Operator

OpenStudy (perl):

openstudy says errors

OpenStudy (jhannybean):

Here, @perl http://prntscr.com/4wnm9o

OpenStudy (anonymous):

one example

OpenStudy (anonymous):

http://prntscr.com/4wnncd

OpenStudy (anonymous):

@amistre64 pls. help me ..

OpenStudy (perl):

its better if you rewrite your original expression

OpenStudy (kainui):

These are linear operators, and they don't commute in general. For example, does multiplying by x and taking the derivative as two operators commute? Well, let's try to do them on some random function y and see what happens: First, the operators x with a hat just means multiply by x. D with a hat means take the derivative. \[\LARGE \hat x (\hat D(y))=\hat x \frac{dy}{dx}=x*\frac{dy}{dx}\] Now let's reverse the order \[\LARGE \hat D ( \hat x(y))= \hat D(x *y)=1*y+x \frac{dy}{dx}\] Since \[\LARGE x y' \ne y + x y'\] then we say that these operators don't commute since order of operation mattes! This is one of the main rules that makes operator algebra different than normal algebra. So now when you look at G and H you have to be careful when you multiply them together than you keep the multiplication straight, otherwise it will be wrong. You should expect GH to not be equal to HG, so if you get the same answer for both you might be doing something wrong. Give it your best shot, and I'll help you finish it. =)

OpenStudy (kainui):

One last note, if I slightly change how I wrote it so it looks more like what you were give it might be more clear to you. So if I strip away all the stuff we were talking about and just leave the operators, it will look like this: \[\LARGE xD \ne Dx\]

OpenStudy (anonymous):

i'll post my sol'n later... I hope you'll help me understand this.. thanks a lot!!!!!! later..

OpenStudy (anonymous):

\[G=D(xD-1) \]\[H=(xD-1)(xD+2)\] \[GHy=D(xD-1)[(xD-1)(xD+2)(y)]\] I AM STOCK,.. how can I multiply (xD-1)(xD+2)(y)??

OpenStudy (anonymous):

@Kainui

OpenStudy (kainui):

Just distribute like you normally would. As an example: (A+B)(A-B)= A^2+BA-AB-B^2 The middle two terms don't cancel because we can't say BA=AB like we usually do in commutative algebra.

OpenStudy (anonymous):

\[x^2 +2xD - xD - 2\] ?

OpenStudy (anonymous):

\[xD(xD+2) - xD - 2 \] ?? @ Kainui

OpenStudy (anonymous):

@Kainui

OpenStudy (anonymous):

@amistre64 ?

OpenStudy (amistre64):

im not familiar with the D notation, ive seen it around, but i just dont have the practice needed to make it set in my head.

OpenStudy (amistre64):

Dy means, dy/dx ? or something else

OpenStudy (amistre64):

got a sideways picture lol yeah, Dy means dy/dx

OpenStudy (amistre64):

G = D(xD-1) H = (xD-1)(xD+2) Determine GH and HG. is that G(Hy) and H(Gy)

OpenStudy (anonymous):

yes.. :)

myininaya (myininaya):

I wonder xD-1 means (or xD+2) I know D means d/dx but I don't understand what are we taken the derivative of since there is nothing to follow the D.

OpenStudy (amistre64):

Gy = Dy (x Dy-1) Hy = (x Dy-1)(x Dy+2) Dy = y' sooo Gy = y' (x y'-1) Hy = (xy'-1)(xy'+2) ---------------------- G(Hy) = D (x D-1) (xy'-1)(xy'+2) G(Hy) = (xD^2 - D) (xy'-1)(xy'+2) G(Hy) = xD^2[(xy'-1)(xy'+2)] - D[(xy'-1)(xy'+2)]

OpenStudy (amistre64):

G(Hy) is essentially: x times the second derivative of Hy, minus the first derivatrive of Hy

OpenStudy (amistre64):

\[D[(xy'-1)(xy'+2)] \] \[(xy'-1)'(xy'+2) + (xy'-1)(xy'+2)'\] \[(x'y'+xy''-1) (xy'+2) + (xy'-1)(x'y'+xy''+2)\] and the second derivative is just one more after that

OpenStudy (amistre64):

pfft, -1 to 0 and +2 to 0

OpenStudy (amistre64):

\[(x'y'+xy'') (xy'+2) + (xy'-1)(x'y'+xy'')\] \[(y'+xy'') [(xy'+2) + (xy'-1)]\] \[D(Hy)=(y'+xy'') (2xy'+1)\]

OpenStudy (amistre64):

so, does this make sense?

OpenStudy (anonymous):

why we must derive Hy?

OpenStudy (amistre64):

G(Hy) is essentially: x times the second derivative of Hy, minus the first derivatrive of Hy

OpenStudy (anonymous):

use Dy instead of y' pls? im confused......

OpenStudy (amistre64):

youll have to do the modifications, im using what makes the most sense to me ...

OpenStudy (anonymous):

okay.. ill try... sorry :)

OpenStudy (amistre64):

just change the y' to Dy, y'' to D^2y etc ... but as far as processing it, im gonna have to use the prime notation to make sure i dont mess up

OpenStudy (anonymous):

Gy= Dy(xDy-1) why not Gy= Dy(xDy-y) ? :)

OpenStudy (amistre64):

hmmm, let me redo that ..... G = D(xD-1) G = xD^2 - D Gy = xD^2y - Dy Gy = D(xDy - y) D is not a variable, its an operator. we need to treat it like an operator then. I most likely misthought Gy to start with ... lets chk up my Hy :)

OpenStudy (amistre64):

H = (xD-1)(xD+2) H = x^2 D^2 +xD -2 Hy = x^2 D^2y +xDy -2y Hy = xD(x Dy +y) -2y

OpenStudy (amistre64):

Gy = D(xDy - y) does this simplify to: GHy = D(xDHy - Hy) ??

OpenStudy (anonymous):

D(xD-1) product rule? =(xD^2 + D - D) =xD^2 ?

OpenStudy (amistre64):

\[D(xD-1)\] \[\frac d{dx}(x\frac d{dx}-1)\] \[x\frac {d^2}{dx^2}-\frac d{dx}\]

OpenStudy (amistre64):

assuming of course that im catching on to this stuff ... correct me as you see fit

OpenStudy (amistre64):

explain why you want to use a product rule

OpenStudy (anonymous):

x'D + xD' ?

OpenStudy (amistre64):

d/dx is not a variable (or a function), its an operator. it only has a definition when it actually operates on something d/dx of y = dy/dx d/dx of x^2 = 2x dx/dx = 2x d/dx of ax^2 = da/dx x^2 + a 2x dx/dx i see why you are wanting to apply it to x * d/dx but im just not sure if thats applicable. can you show me a rule or thrm for it?

OpenStudy (anonymous):

uhhmm that's what my prof. taught us..... :/

OpenStudy (amistre64):

D (xD-1) y D(x Dy - y) Dx Dy + xD^2y - Dy Dx = 1 assuming D = d/dx Dx Dy + xD^2y - Dy Dy + xD^2y - Dy xD^2y = xy'' if we use the y to define the operator as a function, then i can see it as that

OpenStudy (amistre64):

(xD-1) Dy xD^2y -Dy = xy'' - y' so commutability is not gaurenteed with out constant stuff to me

OpenStudy (amistre64):

D(xy' - y) (x'y + xy'' - y') yeah, im going to need alot more practice at this then i currently have to be of any use in these kinds of things :/

OpenStudy (anonymous):

:) me too haha it is late already... need to sleep BIG THANKS TO YOU!!! hope I can finish answering this tomorrow.. night! :)

OpenStudy (amistre64):

good luck, if i can review this today i will and if a come up with anything profound .... ill try to let you know :)

OpenStudy (kainui):

Nonononono we are not taking any derivatives, we are purely doing operator algebra! We are never once taking an actual derivative. The only reason I showed derivatives was to justify that operator algebra is not commutative. Everything else is basically the same, and if you decide to, you can operate on functions with them later. So I guess I'll post the answer since this has gone on too long.

OpenStudy (kainui):

If G = D(xD-1) and H = (xD-1)(xD+2) Determine GH and HG. So let's distribute out G and H separately \[\large G=D(xD-1)=D(xD)-D=D+xD^2-D=xD^2\] See how I did the "product rule" here? We can test it by running a function through it to make sure: \[\large D(xD-1)y=D(xy'-y)=y'+xy''-y'=xy''=xD^2y\] In fact, we can use "running a function through" as our method of simplifying it and just "factor out" the differentiation operator like I did in the last step. Remember when I said earlier that \[\LARGE xD \ne Dx\] you should be incredibly cautious and go up and reread my explanation above, as this is key, especially since xD shows up here. So now, we need to simplify H. I'll save that for the next post.

OpenStudy (kainui):

Ok you should have attempted the next one on your own so that you can make sure you're doing it correctly. This is also for @amistre64 so you can learn too! =D \[\large H=(xD-1)(xD+2)=xD(xD)+xD2-xD-2\] since differentiation is a linear operator, multiplying by a scalar actually commutes with everything. So you can safely say xD2-xD=xD. Just as a quick reminder that we're talking about derivatives, here's an example of this: \[\LARGE \frac{d}{dx}(3x^2)=3 \frac{d}{dx}(x^2)=6x\] See how we can slide scalars around without consequences? Coolio. Now notice we have the same D(xD) we had last time, but now we have an x in there. This is just multiplying afterwards, still nothing crazy going on, it's only derivatives that should give you some hesitation. So now we have as our final result for H: \[\LARGE H=x^2D^2+2xD-2\] Now we can multiply them together the same route we've been going. Like I said earlier, you can just take an arbitrary function y, find Hy, then find GHy and after you have simplified it factor out the D's from the y's and throw the y's away to get the GH operator. Rinse and repeat for HG.

OpenStudy (anonymous):

\[G=xD^2 \] \[Hy=x^2D^2y+2xDy-2y \] \[GHy= xD^2 (x^2D^2y+2xDy-2y)\] \[=xD(x^2D^3y+2xD^2y+2xD^2y+2Dy-2Dy)\] \[=xD(x^2D^3y+4xD^2y)\] \[=x(x^2D^4y+2xD^3y+4xD^3y+4D^2y)\] \[GHy=x^3D^4y+6x^2D^3y+4xD^2y \] \[GH=x^3D^4+6x^2D^3+4xD^2\] ???

OpenStudy (anonymous):

\[HGy=(x^2D^2+2xD-2)(xD^2y)\] \[=x^2D^2(xD^2y) + 2xD(xD^2y)-2xD^2y\] \[=x^2D(xD^3y+D^2y)+2x(xD^3y+D^2y)-2xD^2y\] \[=x^2(xD^4y+D^3y+D^3y)+2x^2D^3y+2xD^2y-2xD^2y\] \[=x^3D^4y+2x^2D^3y+2x^2D^3y \] \[HGy=x^3D^4y+4x^2D^3y \] \[HG=x^3D^4+4x^2D^3\] ???

OpenStudy (anonymous):

did i do it right? @Kainui

OpenStudy (anonymous):

@abdullah1995 hi can you help me? :)

OpenStudy (abdullah1995):

I am sorry but either i don't understand what's going on or the math is too advanced for me :c

OpenStudy (anonymous):

it's alright.. :) hope you'll learn from this too :D

OpenStudy (abdullah1995):

indeed. what university year is this question ?

OpenStudy (abdullah1995):

damn lol. i just started university like 2 months ago for electrical and computer engineering

OpenStudy (anonymous):

@Kainui ??

OpenStudy (anonymous):

hmm... THANK YOU VERY MUCH @Kainui !!!